Hanson Inc. has the following material standard to manufacture one unit of product:

1.5 pounds per Unit at $4.00 per pound
Last week, 2,800 pounds of material were purchased at a total cost of $10,920, and 1,700 pounds were used to make 1,000 units.

3) Hanson’s material price variance (MPV) for the week was:

  1. $280 unfavorable.

  2. $280 favorable.

  3. $800 unfavorable.

  4. $800 favorable.

4) Hanson’s material quantity variance (MQV) for the week was:

  1. $5200 unfavorable.

  2. $5200 favorable.

  3. $800 unfavorable.

  4. $800 favorable.
Answer #1

3) Material price variance = (Standard price-actual price)actual qty = (4*2800-10920) = 280 F

So answer is b) $280 Favorable

4) Material quantity variance = (Standard qty-actual qty)Standard price = (1000*1.5-1700)*4 = 800 U

So answer is c) $800 Unfavorable